OpenStudy (anonymous):

Solve the equation below for the variable (y): 2x + 5y = 20 A)y = 5times(20 - 2x) B)y=2x-20/5 C)y=18x/5 D)y=20-2x/5

OpenStudy (igreen):

\(2x + 5y = 20\) Subtract 2x to both sides: \(5y = 20 - 2x\) Divide 5 to both sides: \(y = \dfrac{20 - 2x}{5}\)

OpenStudy (igreen):

Therefore, your answer is D.

OpenStudy (anonymous):

Thanks

OpenStudy (igreen):

No problem.